OpenStudy (amonoconnor):

When finding the anti-derivative of an expression, if it's a fraction, does one have to use the Quotient Rule? *(We started anti derivatives today in Calc... With a long-term substitute, who is a math major, but not a teacher by degree, at all. Thus, our notes are incomplete, but assignments still due daily!?!) Specifically, I'm trying to find the anti-derivative of: "-2/x^3" Any and all help is greatly appreciated! :)

OpenStudy (amonoconnor):

Basically... I know you have to add one to the exponent of a term, then divide by the new exponent, and add C. I know that you address terms individually in polynomials, but this (above) looks like one term to me... So... Do I ignore the top until I divide the bottom? Meh! I am clueless!

ganeshie8 (ganeshie8):

rewrite \(\large \frac{-2}{x^3}\) as \(-2x^{-3}\)

OpenStudy (amonoconnor):

And then proceed, well, normally?

ganeshie8 (ganeshie8):

yes, then use below : \(\large \int x^n = \frac{x^{n+1}}{n+1} + c\)

OpenStudy (amonoconnor):

This guy teaching Calc to us told me to "not change the sign of the exponents in the problem by rearranging". Is that valid?

OpenStudy (isaiah.feynman):

Well @amonoconnor you integrate the function. Use the form @ganeshie8 showed

OpenStudy (isaiah.feynman):